# Not All Clients Are Equal: Personalized Federated Learning on Heterogeneous Multi-Modal Clients

### v1

@@ 4.1 Similarity-Aware Aggregation
where $\mathcal{D}_{i}$ is the data stream of $i_{\text{th}}$ client seen up to this point, $\ell$ refers to the loss function, $z\subset\mathcal{D}_{i}$ is the mini-batch in the data stream, and $W_{s}$ is a small-scale frozen pre-trained model $W_{s}$. For computational efficiency, we: (1) use gradients from a small-scale frozen pre-trained model $W_{s}$ that provides sufficient representativeness with reduced overhead (lee2024concept, ; maharana2025adaptinfty, ), and (2) compute only the last-layer gradient, as preceding layers’ gradients are proportional to it based on the chain rule (koh2023online, ; seo2025budgeted, ). Moreover, we compute the gradient $g_{i}$ in every $m$ iterations rather than in every batch, thus incurring negligible additional cost, as shown in Sec. A.6. Note that we measure gradients of a frozen pre-trained model, not the training model. This is because in heterogeneous PFL, clients train on diverse tasks, such as VQA and visual reasoning, and gradient similarities from models trained on heterogeneous data may not capture actual similarity (tang2020similarity, ; evans2024data, ).
Decayed Client-Wise Gradient $\hat{g}_{i}$.
However, $g_{i}$ may fail to reflect task similarity under shifting data distributions, as it is an expectation over the entire data stream $D_{i}$ (Eq. 3), i.e., unweighted average of gradients across all time points, ignoring forgetting of the model over time. Consequently, client 1 learning $A\rightarrow B$ and client 2 learning $B\rightarrow A$ yield the same $g_{i}$, despite retaining different knowledge due to catastrophic forgetting (McCloskey1989catastrophic, ; ratcliff1990connectionist, ). To reflect the model knowledge shifts under distribution shifts, we introduce decayed gradient similarity $\hat{g}_{i}$, computed using exponential moving average (EMA) of past gradients, inspired by the exponential decay of knowledge in forgetting (mahto2021multi, ; chien2021slower, ; seo2025budgeted, ). Formally, for the $i_{\text{th}}$ client, $\hat{g}_{i}(t)$ with EMA ratio $\alpha$ at timestep $t$ is defined recursively as:
$\hat{g}_{i}(t)=(1-\alpha)*\hat{g_{i}}(t-1)+\alpha*g_{i}(t),$ (4)
@@
Using the decayed client-wise gradients $\{\hat{g}_{1},\dots,\hat{g}_{N}\}$ from $N$ clients, we construct a client-similarity matrix $S\in\mathbb{R}^{N\times N}$, where $S_{ij}=\text{cos}(\hat{g}_{i},\hat{g}_{j})$. The customized global module for the $i_{\text{th}}$ client, $G_{i}$, is then constructed by weighted aggregation of local modules =$\mathcal{L}=\{L_{1},\dots,L_{N}\}$ as follows:
$G_{i}=\sum_{j=1}^{N}w_{ij}~{}L_{j},~{}~{}~{}~{}w_{ij}=\frac{e^{{\text{cos}(\hat{g}_{i},\hat{g}_{j})}/\tau}}{\sum_{n=1}^{N}e^{\text{cos}(\hat{g}_{i},\hat{g}_{n})/\tau}},$ (5)
where $\tau$ denotes the softmax temperature and cos denotes the cosine similarity.
@@ 4.2 PQ-LoRA
LoRA matrices $A\in\mathbb{R}^{r\times d_{I}}$ and $B\in\mathbb{R}^{d_{O}\times r}$ depend on model-specific hidden dimensions $d_{I}$ and $d_{O}$, and thus cannot be directly aggregated across different models. To share knowledge among heterogeneous architectures, we introduce PQ-LoRA, which inserts dimension-invariant modules $P\in\mathbb{R}^{r\times r}$ and $Q\in\mathbb{R}^{r}$ between $A$ and $B$. The dimensions of these modules depend only on the low-rank size $r$, making them shareable across heterogeneous architectures. We illustrate a comparison of PQ-LoRA with conventional LoRA in Fig. 3. Formally, PQ-LoRA outputs $h_{O}$ as:
@@
Figure 3: The illustration of (a) Conventional LoRA and (b) PQ-LoRA. While $A$ and $B$ are trainable in conventional LoRA, PQ-LoRA freezes both, updating only the dimension-invariant modules $P\in\mathbb{R}^{r\times r}$ and $Q\in\mathbb{R}^{r}$ during training.
Although $P$ and $Q$ are shareable across heterogeneous architectures having different dimensions, two main challenges hinder their aggregation: (i) depth heterogeneity - heterogeneous architectures often differ in depth, making it non-trivial to determine which layers across models should be aggregated; and (ii) weight misalignment - interpolating weights with different optimization trajectories can degrade performance (entezari2021role, ; jordanr2023epair, ; neyshabur2020being, ; stoica2025model, ). To address these challenges, we adopt two strategies: (i) block-wise aggregation, which aligns layers at the same relative depth, and (ii) weight alignment, ensuring heterogeneous models share the same initialization. For simplicity, we consider aggregating two heterogeneous models, though it naturally extends to multiple heterogeneous models.
@@ 4.2.1 Block-Wise Aggregation
We measure layer-wise representation alignment between two MLLMs, $W_{i}$ and $W_{j}$, with different depths using CKA (kornblith2019similarity, ), and observe high similarity between layers at similar relative depths, i.e., approximately linear alignment (Sec. A.7). Accordingly, we divide each model into $N_{B}$ blocks linearly and attach a PQ-LoRA to each block’s final layer, enabling cross-model sharing at relevant depths. $I_{k}$, the attachment layer index of the $k$-th PQ-LoRA in a $|W|$-layer model, is defined as:
@@
4.2.2 Weight Alignment of PQ-LoRA
@@
Motivated by findings that models fine-tuned from the same initialization share optimization paths and can be merged without interference (wortsman2022model, ; wortsman2022robust, ; yadav2023ties, ), we set the pair $X$ to share a common initialization. Dimension-invariant modules, $P_{i},P_{j}\in\mathbb{R}^{r\times r}$ and $Q_{i},Q_{j}\in\mathbb{R}^{r}$, can share the same weight directly, while dimension-dependent modules, $A_{i}\in\mathbb{R}^{r\times d_{i}}$, $A_{j}\in\mathbb{R}^{r\times d_{j}}$ and $B_{i}\in\mathbb{R}^{d_{i}^{\prime}\times r}$, $B_{j}\in\mathbb{R}^{d_{j}^{\prime}\times r}$, cannot due to dimension mismatches. Although $P$ and $Q$ are only shared between heterogeneous models $M_{i}$ and $M_{j}$, $A$ and $B$ should also be aligned, as they affect the output $h_{O}$ (Eq. 6), thereby affecting optimization trajectories. Therefore, we align $A_{i},A_{j}$ and $B_{i},B_{j}$, ensuring aligned representations for the same input, even in parameter spaces with different dimensions.
Aligning $A$ Matrices.
@@
We first align $A_{i}\in\mathbb{R}^{r\times d_{i}}$ and $A_{j}\in\mathbb{R}^{r\times d_{j}}$ using MSE loss in the shared $r$-dimensional space with publicly available data $\mathcal{D}_{p}$. For $\mathcal{D}_{p}$, we use a subset of pre-training data of the employed MLLMs, rather than any downstream client data. In particular, we use the smaller model $A_{i}$ (where $d_{i}<d_{j}$) as the pivot model by freezing it and updating the larger model $A_{j}$ to minimize the L2 loss on $r$-dimensional representations obtained from public data $\mathcal{D}_{p}$, as:
$\min_{A_{j}}\frac{1}{|\mathcal{D}_{p}|}\sum_{(x,y)\in\mathcal{D}_{p}}\text{MSE}(A_{i}(x),A_{j}(x)),$ (8)
where $A(x)$ denotes $r$-dimensional feature extracted by $A$ for input $x$.
Aligning $B$ Matrices.
@@
We then align $B_{i}$ and $B_{j}$. Since they have different output dimensions $d^{\prime}_{i}\neq d^{\prime}_{j}$, MSE loss is inapplicable. Instead, we employ canonical correlation analysis (CCA) (DeBie2005, ), which finds projection matrices that maximize the correlation between two feature sets. Using $m$ public data $\{(x_{k},y_{k})\}_{k=1}^{m}\subset D_{p}$, we extract output features $H_{i}\in\mathbb{R}^{m\times d^{\prime}_{i}}$ and $H_{j}\in\mathbb{R}^{m\times d^{\prime}_{j}}$ from $B_{1}$ and $B_{2}$, then find projection matrices $\Pi_{i}$ and $\Pi_{j}$, which project them to the maximally correlated space. We project $B_{i}$ to the shared maximally correlated space, i.e., $(\Pi_{i})^{T}\cdot B_{i}$, and subsequently bring it to $B_{j}$ space by inverse projection, as $B_{j}=({\Pi_{j}}^{-1})^{T}\cdot(\Pi_{i})^{T}\cdot B_{i}$.
Further Details.
@@
After aligning PQ-LoRA, we update shareable modules $P$ and $Q$, while freezing $A$ and $B$ during local training to preserve alignment. We provide theoretical justification for this freezing approach in Sec. A.1. Moreover, during alignment, we enforce orthogonality in $A$ and $B$ to maximize the expressive capacity (i.e., span) of PQ-LoRA weight updates, following Theorem 1. We provide details of enforcing orthogonality in Sec. A.4 and the proof of Theorem 1 in Sec. A.2.
@@
After receiving the aggregated PQ-LoRA (i.e., global PQ-LoRA) by SIMA at each communication round, clients freeze it during training to preserve global knowledge and update only the local model for personalization. Specifically, at the $l_{\text{th}}$ layer, given an input hidden state $h_{I}$, we combine the output of the local LoRA (i.e., $h_{L}$), the frozen global LoRA (i.e.,$h_{G}$), and the pre-trained weights (i.e., $W_{p}h_{I}$), by adaptively balancing them using a learnable gating parameter $\beta$. With sigmoid-normalized balancing parameter $\tilde{\beta}=\sigma(\beta)$, the output hidden state $h_{O}$ is computed as follows:
@@ 5 DRAKE Benchmark
We propose a comprehensive multi-modal FL benchmark, called DRAKE, with three key merits: (i) Task heterogeneity: Each client handles distinct multi-modal tasks (e.g., visual reasoning, VQA), while existing benchmarks merely assign non-i.i.d. subsets split from a single dataset. (ii) Dynamic distribution: Client datasets contain progressive tasks (e.g., encountering new visual concepts), simulating real-world distribution shifts. To our knowledge, DRAKE is the first benchmark that supports the multi-modal federated continual learning scenario. (iii) Evaluation for generalizability: DRAKE incorporates unseen task data to evaluate the generalization and fast adaptation capabilities of FL models. We summarize the comparison with existing FL benchmarks in Tab. 1.
As shown in Fig. 4, DRAKE consists of three training task subgroups, i.e., VQA, visual relation, and multi-modal reasoning, and two unseen task subgroups, i.e., new task types and new visual domains. Specifically, it comprises 40 heterogeneous tasks sourced from 19 different multi-modal datasets, totaling 375k images and 274k questions. We provide more details of DRAKE in Sec. A.18.
Dataset Multi-Image Support Multi-Data Sources Distribution Shifts Multi- Modalities Unseen Evaluation
@@
Table 1: Comparison of FL benchmarks across key dimensions: Multi-Image Support (handling multiple images per input), Multi-Data Sources (using diverse datasets vs. non-i.i.d. splits of single dataset), Distribution Shifts (evolving client data distributions), Multi-Modalities (single or multiple data types), and Unseen Evaluation (testing on tasks unseen during training).

### v3

@@ 4.1 Relevance-Guided Aggregation
where $\mathcal{D}_{i}$ is $i_{\text{th}}$ client’s data stream, $\ell$ refers to the loss function, $z\subset\mathcal{D}_{i}$ is a mini-batch, and $W_{s}$ is a small-scale frozen pre-trained model $W_{s}$. For efficiency, we (i) use gradients from a small-scale frozen pre-trained model $W_{s}$ that provides sufficient representativeness with reduced overhead (Lee et al., 2024), and (ii) compute only the last-layer gradient, as preceding layers’ gradients are proportional to it based on the chain rule (Seo et al., 2025). Moreover, we compute the gradient $g_{i}$ every $m$ batch iterations rather than every batch, thus incurring negligible additional cost, as shown in Sec. A.8. Note that we measure gradients from a frozen pre-trained model, not the actual training model. This is because in heterogeneous PFL, clients train on diverse tasks, and gradient similarities from models trained on heterogeneous data may not capture task similarity (Tang et al., 2020; Evans et al., 2024).
Decayed Client-Wise Gradient $\hat{g}_{i}$. However, $g_{i}$ may not reflect learned task relevance under shifting data distributions, as it is an expectation over the entire data stream $D_{i}$ (Eq. 1), i.e., unweighted average of gradients across all time points, ignoring forgetting of the model over time. Consequently, client 1 learning $A\rightarrow B$ and client 2 learning $B\rightarrow A$ yield the same $g_{i}$, despite retaining different knowledge due to catastrophic forgetting (McCloskey & Cohen, 1989; Ratcliff, 1990). To reflect the model knowledge shifts under distribution shifts, we introduce decayed gradient $\hat{g}_{i}$, computed using the exponential moving average (EMA) of past gradients, inspired by the exponential decay of knowledge in forgetting (Mahto et al., 2021; Chien et al., 2021; Seo et al., 2025). Formally, for the $i_{\text{th}}$ client, $\hat{g}_{i}(t)$ with EMA ratio $\alpha$ at timestep $t$ is defined recursively as:
$\hat{g}_{i}(t)=(1-\alpha)\cdot\hat{g_{i}}(t-1)+\alpha\cdot g_{i}(t),$ (2)
@@
Sanitized Client-Wise Gradient $\tilde{g}_{i}$. Transmitting EMA-aggregated gradients (i.e., $\hat{g}_{i}$) rather than per-sample gradients mitigates gradient-based privacy attacks, as gradient mixing (i.e., aggregating) increases resistance to gradient inversion (Mo et al., 2021). We further prevent privacy risks by transmitting sanitized gradient $\tilde{g}_{i}$ through: (i) adding Gaussian noise $\epsilon$ to $\hat{g}_{i}$ and (ii) applying gradient compression (i.e., randomly selecting only $N_{s}\%$ of the gradient vector dimensions from $\hat{g_{i}}\in\mathbb{R}^{d}$), as randomly sampled dimensions can approximate full-gradient distributions (Li et al., 2023b) while making gradient inversion substantially more difficult than using full gradients (Zhu et al., 2019) and simultaneously reducing transmission costs (detailed in Sec. A.9). Formally, $\tilde{g}_{i}$ is defined as:
$\tilde{g}_{i}(t)=\mathbf{M}\odot(\hat{g_{i}}(t)+\mu\epsilon),\quad\epsilon\sim\mathcal{N}(0,\bm{I}_{d})$ (3)
where $\mathbf{M}\in\{0,1\}^{d}$ is the binary mask for random subsampling and $\mu$ denotes the noise scale.
Using the sanitized client-wise gradients $\{\tilde{g}_{1},\dots,\tilde{g}_{N}\}$ from $N$ clients, we construct a client-relevence matrix $S\in\mathbb{R}^{N\times N}$, where $S_{ij}=\text{cos}(\tilde{g}_{i},\tilde{g}_{j})$. The customized global module for the $i_{\text{th}}$ client, $G_{i}$, is then constructed by weighted aggregation of local modules $\mathcal{L}=\{L_{1},\dots,L_{N}\}$ as follows:
$G_{i}=\sum_{j=1}^{N}w_{ij}~L_{j},~~~~w_{ij}=\frac{e^{{\text{cos}(\tilde{g}_{i},\tilde{g}_{j})}/\tau}}{\sum_{n=1}^{N}e^{\text{cos}(\tilde{g}_{i},\tilde{g}_{n})/\tau}},$ (4)
where $\tau$ denotes the softmax temperature and $\text{cos}(\cdot,\cdot)$ denotes cosine similarity.
@@ 4.2 PQ-LoRA
Figure 3: Illustration of (a) Conventional LoRA and (b) PQ-LoRA. While AA and $B$ are trainable in conventional LoRA, PQ-LoRA freezes both, updating only the dimension-invariant modules $P\in\mathbb{R}^{r\times r}$ and $Q\in\mathbb{R}^{r}$ during training.
LoRA matrices $A\in\mathbb{R}^{r\times d_{I}}$ and $B\in\mathbb{R}^{d_{O}\times r}$ depend on model-specific hidden dimensions $d_{I}$ and $d_{O}$, preventing direct aggregation across different models. To enable knowledge sharing among heterogeneous architectures, we introduce PQ-LoRA, which inserts dimension-invariant modules $P\in\mathbb{R}^{r\times r}$ and $Q\in\mathbb{R}^{r}$ between $A$ and $B$. Their dimensions depend only on the low-rank $r$, making them shareable across heterogeneous models. We illustrate a comparison of PQ-LoRA with conventional LoRA in Fig. 3. Formally, PQ-LoRA outputs $h_{O}$ as:
@@
Although $P$ and $Q$ are shareable across heterogeneous architectures having different dimensions, two main challenges hinder their aggregation: (i) depth heterogeneity - heterogeneous architectures often differ in depth, making it non-trivial to determine which layers across models should be aggregated; and (ii) weight misalignment - interpolating weights with different optimization trajectories can degrade performance (Jordan et al., 2023; Stoica et al., 2025). We address these challenges with two strategies: (i) block-wise aggregation, which aligns layers at the same relative depth, and (ii) weight alignment, ensuring heterogeneous models share the same initialization. For simplicity, we consider aggregating two heterogeneous models, though it naturally extends to multiple heterogeneous models.
@@ 4.2.1 Block-Wise Aggregation
Figure 4: Layer-wise similarity between Llama-1B and Llama-3B measured with CKA. The diagonal brightest band shows the strongest alignment between layers at similar relative depths (e.g., Llama-1B layer 8 - Llama-3B layer 14).
To decide which layers of heterogeneous models should share knowledge, we measure layer-wise representation alignment between two MLLMs, $W_{i}$ and $W_{j}$, of different depths using CKA (Kornblith et al., 2019). We observe high similarity between layers at similar relative depths, i.e., approximately linear alignment (Fig. 4). Accordingly, we divide each model into $N_{B}$ blocks linearly and attach a PQ-LoRA to each block’s final layer, enabling cross-model sharing at relevant depths. The attachment layer index $I_{k}$ of the $k$-th PQ-LoRA in a $|W|$-layer model is defined as:
@@
4.2.2 Weight Alignment in PQ-LoRA
@@
Motivated by findings that models fine-tuned from the same initialization share optimization paths and can be merged without interference (Wortsman et al., 2022a; b; Yadav et al., 2023), we set the pair $X$ to share a common initialization. Dimension-invariant modules, $P_{i},P_{j}\in\mathbb{R}^{r\times r}$ and $Q_{i},Q_{j}\in\mathbb{R}^{r}$, can share the same weight directly, but dimension-dependent modules, $A_{i}\in\mathbb{R}^{r\times d_{I}^{(i)}}$, $A_{j}\in\mathbb{R}^{r\times d_{I}^{(j)}}$ and $B_{i}\in\mathbb{R}^{d_{O}^{(i)}\times r}$, $B_{j}\in\mathbb{R}^{d_{O}^{(j)}\times r}$, cannot due to dimension mismatches. Although only $P$ and $Q$ are shared between heterogeneous models $M_{i}$ and $M_{j}$, $A$ and $B$ should also be aligned, as they affect the output $h_{O}$ (Eq. 5) and optimization trajectories. Therefore, we align $A_{i},A_{j}$ and $B_{i},B_{j}$, ensuring aligned representations for the same input, even in parameter spaces with different dimensions.
@@
Aligning $A$ Matrices. We first align $A_{i}\in\mathbb{R}^{r\times d_{I}^{(i)}}$ and $A_{j}\in\mathbb{R}^{r\times d_{I}^{(j)}}$ using L2 loss in the shared $r$-dimensional space with publicly available data $\mathcal{D}_{p}$. For $\mathcal{D}_{p}$, we use a subset of the MLLMs’ pre-training data, as detailed in Sec. A.17. In particular, we use the smaller model $A_{i}$ (where $d_{I}^{(i)}<d_{I}^{(j)}$) as the pivot model by freezing it and updating the larger model $A_{j}$ to minimize the L2 loss on $r$-dimensional representations obtained from $\mathcal{D}_{p}$, as $\min_{A_{j}}\frac{1}{|\mathcal{D}_{p}|}\sum_{(x,y)\in\mathcal{D}_{p}}\|(A_{i}(x)-A_{j}(x)\|^{2}_{2}$, where $A(x)$ denotes $r$-dimensional feature extracted by $A$ for input $x$.
@@
Aligning $B$ Matrices. We then align $B_{i}$ and $B_{j}$. Since their output dimensions differ ($d_{O}^{(i)}\neq d_{O}^{(j)}$), L2 loss is inapplicable. Instead, we employ canonical correlation analysis (CCA) (De Bie et al., 2005), which finds projection matrices maximizing the correlation between two feature sets. Using $m$ public data $\{(x_{k},y_{k})\}_{k=1}^{m}\subset D_{p}$, we extract output features $H_{i}\in\mathbb{R}^{m\times d_{O}^{(i)}}$ and $H_{j}\in\mathbb{R}^{m\times d_{O}^{(j)}}$ from $B_{i}$ and $B_{j}$, then find projection matrices $\Pi_{i}\in\mathbb{R}^{d_{O}^{(i)}\times r}$ and $\Pi_{j}\in\mathbb{R}^{{d_{O}^{(j)}\times r}}$, which project them to the maximally correlated space of dimension $r$ (i.e., $H_{i}\Pi_{i}\approx H_{j}\Pi_{j}$). We project $B_{i}$ to the shared maximally correlated space, i.e., $(\Pi_{i})^{T}\cdot B_{i}$, and subsequently bring it to $B_{j}$’s space by inverse projection, as $B_{j}=({\Pi_{j}}^{\dagger})^{T}\cdot(\Pi_{i})^{T}\cdot B_{i}$, where $\Pi_{j}$$^{\dagger}$ denotes the pseudo-inverse of $\Pi_{j}$}.
@@
During the alignment of $A$ and $B$, we enforce orthogonality in $A$ and $B$ to maximize the expressive capacity (i.e., span) of PQ-LoRA weight updates, following Theorem 1. We provide details of enforcing orthogonality in Sec. A.5 and the proof of Theorem 1 in Sec. A.2.
@@
In addition, our method does not incur much computational cost since (i) PQ-LoRA aligning is performed once before federated training to establish a shared initialization, (ii) it is required only for heterogeneous model-type pairs (since some clients may share the same architecture). After aligning $A$ and $B$, we update shareable modules $P$ and $Q$, while freezing $A$ and $B$ during local training to preserve alignment. We provide theoretical justification for this freezing approach in Sec. A.1. This freezing design also reduces the communication cost of PQ-LoRA relative to conventional LoRA, as it requires communication of only $P,Q$ modules. See Sec. A.9 for details of communication costs.
After receiving the aggregated PQ-LoRA (i.e., global PQ-LoRA) by RELA at each communication round, clients freeze it during training to preserve global knowledge and update only the local model for personalization. Specifically, at the $l_{\text{th}}$ layer, given an input hidden state $h_{I}$, we combine the output of the local LoRA (i.e., $h_{L}$), the frozen global LoRA (i.e., $h_{G}$), and the pre-trained weights (i.e., $W_{p}h_{I}$), by adaptively balancing them using a learnable gating parameter $\beta$. With sigmoid-normalized balancing parameter $\tilde{\beta}=\sigma(\beta)$, the output hidden state $h_{O}$ is computed as follows:
@@ 5 DRAKEBenchmark
We propose a novel multi-modal FL benchmark, called DRAKE, with three key merits: (i) Task heterogeneity: Each client handles distinct multi-modal tasks (e.g., visual reasoning, VQA), while existing benchmarks merely assign non-i.i.d. subsets from a single dataset. (ii) Dynamic distribution: Client datasets contain progressive tasks (e.g., encountering new visual concepts), simulating real-world temporal distribution shifts. To the best of our knowledge, DRAKE is the first benchmark supporting multi-modal federated learning under distribution shifts. (iii) Generalizability Evaluation: DRAKE incorporates unseen task data to evaluate the generalizability of FL models. We summarize the comparison with existing FL benchmarks in Tab. 1. DRAKE consists of three training task subgroups, i.e., VQA, visual relation, and multi-modal reasoning, and two unseen task subgroups, comprising 40 heterogeneous tasks sourced from 19 different multi-modal datasets, totaling 375k images and 274k questions, illustrated in Fig. 5. Please refer to Sec. A.24 for more details.
Dataset
Multi-Data
Sources
Distribution
Shifts
Multi-Image
Support
Multi-
Modalities
Unseen
Evaluation
@@
Table 1: Comparison of FL benchmarks across key dimensions: Multi-Data Sources (using diverse datasets vs. non-i.i.d. splits of a single dataset), Distribution Shifts (evolving client data distributions), Multi-Image Support (handling multiple images per input), and Unseen Evaluation (testing on tasks unseen during training). See Sec. A.25 for the detailed comparisons.